OpenStudy (anonymous):

tan²θ-√3tanθ=0. find solutions in the interval [0,360°). Round to the nearest ten minutes. Help!

OpenStudy (anonymous):

\[tan^2(\theta)-\sqrt{3}tan(\theta)=0\] \[tan(\theta)(tan(\theta)-\sqrt{3})=0\] \[tan(\theta)=0\]or \[tan(\theta)=\sqrt{3}\]

OpenStudy (anonymous):

of course you still have to solve for \[\theta\]

OpenStudy (anonymous):

\[tan(\theta)=0\] \[\theta=0\]

OpenStudy (anonymous):

\[tan(\theta)=\sqrt{3}\] \[\theta=\frac{\pi}{3}\] or \[\theta=\frac{4\pi}{3}\] of you ar working in degrees instead of radians it would be 60 and 240